# DynamicFace: High-Quality and Consistent Video Face Swapping using Composable 3D Facial Priors

### v1

@@ Introduction
In this work, to enhance the ability of identity injection, Face Former is proposed to enable the diffusion model to accurately preserve the identity with the face image prompt. In text-to-video tasks, textual prompts primarily ensure semantic relevance with the generated visual content, whereas image-to-video tasks demand precise consistency due to the detailed features encapsulated in images. Prior studies in image-driven generation have used the CLIP image encoder as a substitute for the text encoder in cross-attention layer, but this design struggles with accurate facial representation. Here, we use a face recognition model instead of the vision-language CLIP model (Radford et al. 2021) to extract face embedding from the face image prompt. To effectively inject the face embedding to the original diffusion model, we use a lightweight trainable querying transformer like recent work (Li et al. 2023; Han et al. 2024; He et al. 2024) to transform identity token space to semantic-level textual space. We create $N_{token}$ learnable facial queries as input for the querying transformer. The facial queries will interact with face embedding through each self-attention layer. Thus, these tokens could be sent to the UNet for high-level identity injection in face swapping.
ReferenceNet.
@@
To maintain dense facial textures and preserve detailed identity of source face image, we adopt a ReferenceNet for detailed-level identity injection. The facial features from Face Former is often less effective to preserve detailed information. This is partly because the face encoder processes low-resolution face images (112×112), leading to a loss of fine-grained details. Additionally, face encoder emphasizes high-level feature matching, resulting in a deficit of detailed features during the encoding phase. Inspired by methods like MasaCtrl (Cao et al. 2023) and Animate Anyone (Hu 2024), we adopt a ReferenceNet that improves identity preservation given a source face image. ReferenceNet creates a trainable copy of the base UNet of Stable Diffusion to compute condition features in each denoising step, injecting appearance features through a spatial attention mechanism. The spatial attention mechanism maintains the semantic layout of source face while preserving details such as facial texture, and other intra-personal attributes, ensuring appearance consistency. We integrate features from ReferenceNet into the base UNet by replacing the self-attention layer with a spatial-attention layer in middle and upsample stage, facilitating effective feature learning and initialization. We adopt random cutout on the source face image with irregular shape to enhance the robustness of appearance encoding.
@@
We choose VGGFace2 (Cao et al. 2018), VFHQ (Xie et al. 2022) and a private dataset as the training set. Particularly, low-quality faces are removed to ensure high-quality training. For evaluation, we select FaceForensics++ dataset to test. The training process consists of two stages. In the initial training stage, we sample faces from same individuals to train spatial modules including face querying transformer and ReferenceNet for identity injection, guiders and main denoising UNet for motion controlling. As for the four facial conditions, we randomly dropout either expression-related condition or shape-aware pose condition for their possibly repeated representation to make our network extract information from all conditions, instead of learning from easier one. The purpose of the first stage is to adapt the prior knowledge of Stable Diffusion into human face domain. During the second training stage, we only adjust the temporal layer to maintain the temporal consistency across frames. All the face images and frames are resized to $512\times 512$. We finetune the model for 260,000 steps using a batch size of 32. In the second training stage, we focus on training the temporal layer for 40,000 steps using 16-frame video sequences with a batch size of 16. We use same augmentation in both two stages for robust training. Apart from the reconstruction loss on the whole image from Equation 2, we also add two losses on the whole facial area and several facial regions (e.g., eyes, ears, mouth and nose) to help network focus on the facial appearance and expression. The learning rate is $1\times 10^{-5}$ in two stages. We fine-tuned the 3D face model (Wang et al. 2024b) for higher accuracy. The temporal attention layers are initialized with pretrained model in AnimateDiff (Guo et al. 2024). We optimize the overall framework using Adam on 4 NVIDIA H800 GPUs.
Metrics
@@
Following common practice, we randomly sample 10 images from each video to perform face swapping and compute the image-level metrics including ID retrieval, pose error, expression error, eye motion error and mouth position error. To compute ID Retrieval, we first extract the identity feature using a different face recognition model (Wang et al. 2018) from our face encoder. For each swapped face, we compute the nearest face from all the frames in FF++ using the cosine similarity and check whether it is from the source video. The pose error is computed by the mean L2 distance of pitch, yaw, and roll between the swapped face and the target face estimated by a different pose estimator (Martyniuk et al. 2022). The expression error is the L2 distance between the expression vectors extracted by (Martyniuk et al. 2022) of the swapped face and the target face. Both eye motion error and mouth position error are evaluated by 2D keypoints to compare the position consistency between swapped faces and target faces. We also sample 30 frames to calculate video-level metrics containing CLIP frame consistency and warping error (Lei, Xing, and Chen 2020). For CLIP frame consistency, we use CLIP (Radford et al. 2021) to obtain feature vectors of consecutive frames and then compute the cosine similarity between consecutive frames. To compute warping error, we calculate the warp differences on every two frames with an optical flow estimation network (Ilg et al. 2017) and calculate the final score using the average of all the pairs.
@@
Quantitative Comparisons
Our method is compared with six methods including Deepfakes (DeepFakes 2019), FaceShifter (Li et al. 2020), MegaFS (Zhu et al. 2021), SimSwap (Chen et al. 2020), DiffSwap (Zhao et al. 2023) and Face Adapter (Han et al. 2024). For Deepfakes and FaceShifter, we use their released face swapping results of the sampled 10,000 images. For MegaFS, SimSwap, DiffSwap, and Face-Adapter, the face swapping results are generated with their released codes. Table 1 shows that our method achieves the best scores under most evaluation metrics, including identity retrieval, mouth position error and eye motion error. These results validate the controllable superiority of our method. Our pose error results are slightly poorer than others due to DynamicFace altering face shape, affecting the face shape-sensitive head pose estimator used. Figure 5 illustrates that DynamicFace is competitive in preserving the background, motion of target face, and identity of source face well, especially for shape. In addition, we compare the video face swapping results of DynamicFace with motion module and without motion module to show that even if fine-grained conditions are provided, motion module is still essential due to the strong diversity of diffusion model. Figure 2 shows that our method could achieve promising consistency and generate controllable facial expression across frames. DynamicFace can generate faces with higher quality rather than GAN-based methods, and achieve better controllable ability than other Diffusion-based methods.
Table 2: Ablation study for the motion module of our method during video face swapping.
Identity Preservation Motion Fidelity Temporal Consistency
ID Retri. $\uparrow$ ID Simi. $\uparrow$ Pose $\downarrow$ Expr. $\downarrow$ Mouth $\downarrow$ Eye $\downarrow$ Frame Cons. $\uparrow$ Warping Err. $\downarrow$
A) $w/o$ motion module 99.21 0.594 $1.81$ $3.12$ 1.98 $0.17$ $95.78$ $0.091$
B) $w/$ motion module $98.90$ $0.574$ 1.54 2.94 $2.30$ 0.16 99.02 0.046
Ablation study
In order to further explore the effect of composable face conditions and other modules in DynamicFace, we conduct a ablation study on 1) four facial conditions 2) identity injection modules 3) motion module.
Table 3: Ablation study for the impact of Face Former and ReferenceNet on identity injection. The best result is highlighted in bold.
@@
Significance of composable face conditions.
Four facial conditions aim to provide necessary guidance for controllable generation. Figure 6 shows that each condition plays its role in face swapping and each of them is necessary. Background could provide detailed environmental information, ensuring the generation outside facial area as an inpainting paradigm. Normal map is well-aligned with the shape of source face image and maintains the pose of target face frames, which guide precise shape and pose generation. UV texture map reflects the lighting of target face, making illumination consistent during face swapping.
Reliability of motion module.
@@
Despite our efforts to design precise conditions for generating swapped faces with identical motion to the target face, the intrinsic diversity of diffusion models leads to a noticeable variation. It will generate faces with the same motion but slightly different details, making video face swapping unavailable. Here, we evaluate the importance of motion module from three aspects: identity preservation, motion fidelity, and temporal consistency. Table LABEL:tab:ablation_motion shows that DynamicFace could achieve great temporal consistency with motion module.
Effectiveness of identity injection modules.
ReferenceNet and Face Former intend to enhance the ability of identity preservation with detailed-level and high-level injection. For a fair comparison, we train three models with same settings to explore the importance of identity modules. Experiments in Table 3 indicate that both ReferenceNet and Face Former improve identity extraction. In detail, ReferenceNet could extract detailed textures from the whole reference image and select useful facial features to inject in our main denoising UNet. Face Former provides high-level identity features with the knowledge of face recognition model.
Conclusion
@@
We present a novel method named DynamicFace which leveraged the powerful pretrained diffusion model with delicately disentangled facial conditions, achieving promising face swapping results in both image and video domains. Experiments show that the designed facial conditions could give precise and unique evidence on required information (e.g., shape, expression, pose, lighting, and background). Several efforts have been taken to adapt the diffusion model to face swapping, including Face Former, ReferenceNet, Mixture-of-Guiders and motion module. Extensive experiments demonstrate that our framework can achieve superior results compared to previous methods, with better controllability and scalability. With above fine-grained facial conditions and temporal layer, our method can achieve consistent face swapping in video domain. We hope our attempt can inspire future work to further explore the formulation and implementation of face-swapping to achieve better results.

### v2

@@ 3.4 Appearance-Aware Controlling
Face Former. In this work, to enhance the ability of identity injection, Face Former is proposed to enable the diffusion model to accurately preserve the identity with the face image prompt. In text-to-video tasks, textual prompts primarily ensure semantic relevance with the generated visual content, whereas image-to-video tasks demand precise consistency due to the detailed features encapsulated in images. Prior studies in image-driven generation have used the CLIP image encoder as a substitute for the text encoder in cross-attention layer, but this design struggles with accurate facial representation. Here, we use a face recognition model instead of the vision-language CLIP model [22] to extract face embedding from the face image prompt. To effectively inject the face embedding to the original diffusion model, we use a lightweight trainable querying transformer like recent work [16, 8, 9] to transform identity token space to semantic-level textual space. We create $N_{token}$ learnable facial queries as input for the querying transformer. The facial queries will interact with face embedding through each self-attention layer. Thus, these tokens could be sent to the UNet for high-level identity injection in face swapping.
@@
ReferenceNet. To preserve dense facial textures and detailed identity information from the source face image, we adopt ReferenceNet for fine-grained identity injection. While Face Former encodes low-resolution inputs (112×112), leading to the loss of subtle features due to its emphasis on high-level feature matching, ReferenceNet addresses this limitation by leveraging a trainable replica of Stable Diffusion’s base UNet. Inspired by MasaCtrl [1] and Animate Anyone [12], ReferenceNet integrates appearance features through a spatial attention mechanism during denoising. By replacing self-attention layers with spatial-attention modules in intermediate and upsampling stages, it maintains the semantic layout of the source face while preserving texture details and intra-personal attributes, ensuring robust identity consistency.
@@
3.5 Latent Fusion Total Variation Optimizer
In long-duration video face swapping, abrupt inter-frame motions (e.g., rapid head rotations or expression changes) often lead to temporal flickering and inconsistent identity rendering. To address this, we propose FusionTVO, a plug-and-play module that integrates segment-aware latent fusion and temporal Total Variation (TV) regularization to ensure smooth transitions across frames. While our composable facial conditions and temporal attention layers already ensure coarse temporal alignment, subtle flickering artifacts may persist due to per-frame denoising variations. Given two adjacent segments, overlapping frames near their boundary are assigned fusion weights based on their relative positions. This ensures smooth transitions between segments by prioritizing frames closer to their original segment’s center. To suppress flickering caused by large motions, we impose a TV constraint on consecutive latents without retraining the model. We integrate the TV term into the reverse diffusion process using Half-Quadratic Splitting (HQS) algorithm  [6], decoupling the denoising objective (guided by facial conditions) from the temporal consistency constraint.
$\mathbf{z}_{t-1}^{(1:N)}=\arg\min_{\mathbf{z}}\underbrace{\left\|\mathbf{z}-\hat{\mathbf{z}}_{t-1}^{(1:N)}\right\|_{2}^{2}}_{\text{Denoising Prior }}+\mathcal{R}_{TV}(\mathbf{z}),$ (8)
where $\mathbf{z}_{t-1}^{(1:N)}$ is the initial latent prediction from the diffusion model. For algorithmic details, including pseudocode and hyperparameter, we refer readers to Algorithm 1.
Algorithm 1 FusionTVO
1:Input: Segments{$S_{1},S_{2},...,S_{K}$}, Overlap Size $O$, Facial Conditions $c^{1:N}$, Latent Face Mask $\tilde{M}_{swap}^{1:N}$, Latent Target Video $\mathbf{z}_{\text{target}}^{1:N}$, ${\{\sigma_{t}\}_{t=1}^{T}}$, $\mathbf{z}_{T}^{1:N}$, $\lambda_{\text{fusion}}$, $\lambda_{\text{TV}}$.
2:Output: Fused Face Swapping Video $\mathbf{\hat{z}}_{0}^{1:N}$.
3:for $t=T$ to $1$ do
4:     for $n=1$ to $K$ do $\triangleright$ Denoising steps
5:         $\mathbf{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
6:         $\mathbf{\hat{z}}_{\text{target}}^{1:N}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{\text{target}}^{1:N}+(1-\bar{\alpha}_{t})\mathbf{\epsilon}_{t}$
7:         $\mathbf{\hat{z}}_{0}^{S_{n}}=\frac{1}{\sqrt{\alpha_{t}}}\big{(}\mathbf{{z}}_{t}^{S_{n}}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\epsilon}_{\theta}(\mathbf{{z}}_{t}^{S_{n}},c^{S_{n}},t)\big{)}\circ\tilde{M}_{swap}^{S_{n}}+\mathbf{\hat{z}}_{\text{target}}^{1:N}\circ(1-\tilde{M}_{swap}^{S_{n}})$$\triangleright$ Repaint Background
8:     end for
9:     for each adjacent pair $(S_{k},S_{k+1})$ do
10:         Extract overlapping frames $\{\mathbf{{\tilde{z}}}^{(i)}\}_{i=1}^{O}$
11:         for $i=1$ to $O$ $\triangleright$ Weighted Fusion
12:              $d_{i}=O-i+1,w_{i}=d_{i}\cdot\lambda_{\text{fusion}}$
13:              $\mathbf{{\tilde{z}}}^{(i)}=w_{i}\mathbf{{\hat{z}}}_{S_{k}}^{(i)}+(1-w_{i})\mathbf{{\hat{z}}}_{S_{k+1}}^{(i)}$
14:         end for
15:         Update current overalpping frames in $\mathbf{\hat{z}}_{t-1}^{1:N}$
16:     end for
17:     $\mathbf{\hat{z}}_{t-1}^{1:N}=\mathbf{{\hat{z}}}_{t-1}^{1:N}-{\lambda_{\text{TV}}}\nabla_{\mathbf{\hat{z}}_{t-1}^{1:N}}\|D_{z}\mathbf{\hat{z}}_{t-1}^{1:N}\|^{2}$triangleright TV Opti.
18:end for
19:return $\mathbf{\hat{z}}_{0}^{1:N}$
4 Experiments
4.1 Implementation Details
@@
We choose VGGFace2 [2], VFHQ [30] and a private dataset as the training set. Particularly, low-quality faces are removed to ensure high-quality training. For evaluation, we select FaceForensics++ dataset to test. The training process consists of two stages. In the initial training stage, we sample faces from same individuals to train spatial modules including face querying transformer and ReferenceNet for identity injection, guiders and main denoising UNet for motion controlling. As for the four facial conditions, we randomly dropout either expression-related condition or shape-aware pose condition for their possibly repeated representation to make our network extract information from all conditions, instead of learning from easier one. The purpose of the first stage is to adapt the prior knowledge of Stable Diffusion into human face domain. During the second training stage, we only adjust the temporal layer to maintain the temporal consistency across frames. All the face images and frames are resized to $512\times 512$. We finetune the model for 260,000 steps using a batch size of 32. In the second training stage, we focus on training the temporal layer for 40,000 steps using 16-frame video sequences with a batch size of 16. We use same augmentation in both two stages for robust training. Apart from the reconstruction loss on the whole image from Equation 2, we also add two losses on the whole facial area and several facial regions (e.g., eyes, ears, mouth and nose) to help network focus on the facial appearance and expression. The learning rate is $1\times 10^{-5}$ in two stages. We fine-tuned the 3D face model [29] for higher accuracy. The temporal attention layers are initialized with pretrained model in AnimateDiff [7]. We optimize the overall framework using Adam on 4 NVIDIA H800 GPUs.
4.2 Metrics
@@
Following common practice  [3, 17], we randomly sample 10 images from each video to perform face swapping and compute the image-level metrics including ID retrieval, pose error, expression error, eye motion error and mouth position error. To compute ID Retrieval, we first extract the identity feature using a different face recognition model [26] from our face encoder. For each swapped face, we compute the nearest face from all the frames in FF++ using the cosine similarity and check whether it is from the source video. The pose error is computed by the mean L2 distance of pitch, yaw, and roll between the swapped face and the target face estimated by a different pose estimator [18]. The expression error is the L2 distance between the expression vectors extracted by [18] of the swapped face and the target face. Both eye motion error and mouth position error are evaluated by 2D keypoints to compare the position consistency between swapped faces and target faces. We also sample 30 frames to calculate video-level metrics containing CLIP frame consistency and warping error [15]. For CLIP frame consistency, we use CLIP [22] to obtain feature vectors of consecutive frames and then compute the cosine similarity between consecutive frames. To compute warping error, we calculate the warp differences on every two frames with an optical flow estimation network [13] and calculate the final score using the average of all the pairs.
@@
4.3 Quantitative Comparisons
Our method is compared with six methods including Deepfakes [4], FaceShifter [17], MegaFS [36], SimSwap [3], DiffSwap [34] and Face Adapter [8]. For Deepfakes and FaceShifter, we use their released face swapping results of the sampled 10,000 images. For MegaFS, SimSwap, DiffSwap, and Face-Adapter, the face swapping results are generated with their released codes. Table 1 shows that our method achieves the best scores under most evaluation metrics, including identity retrieval, mouth position error and eye motion error. These results validate the controllable superiority of our method. Our pose error results are slightly poorer than others due to DynamicFace altering face shape, affecting the face shape-sensitive head pose estimator used. Figure 6 illustrates that DynamicFace is competitive in preserving the background, motion of target face, and identity of source face well, especially for shape. Figure 1 shows that our method could achieve promising consistency and generate controllable facial expression across frames. DynamicFace can generate faces with higher quality rather than GAN-based methods, and achieve better controllable ability than other Diffusion-based methods.
4.4 Ablation study
In order to further explore the effect of composable face conditions and other modules in DynamicFace, we conduct a ablation study on 1) four facial conditions 2) identity injection modules 3) temporal consistency module.
Significance of composable face conditions. Four facial conditions aim to provide necessary guidance for controllable generation. Figure 5 shows that each condition plays its role in face swapping and each of them is necessary. Background could provide detailed environmental information, ensuring the generation outside facial area as an inpainting paradigm. Normal map is well-aligned with the shape of source face image and maintains the pose of target face frames, which guide precise shape and pose generation. UV texture map reflects the lighting of target face, making illumination consistent during face swapping. For more detailed ablation studies, please refer to Supplementary Materials.
Identity Preservation Motion Fidelity Video Quality
ID Retri. $\uparrow$ ID Simi. $\uparrow$ Pose $\downarrow$ Expr. $\downarrow$ Mouth $\downarrow$ Eye $\downarrow$ Consistency $\uparrow$ Warp. $\downarrow$ Aesthetic $\uparrow$
A) $w/o.$ motion module 99.21 0.594 $1.81$ $3.12$ 1.98 $0.17$ $95.78$ $0.091$ 3.26
B) $w/.$ motion module $98.90$ $0.574$ 1.54 2.94 $2.30$ 0.16 99.02 0.046 3.48
C) $w/.$ mm & FusionTVO $98.90$ $0.574$ 1.53 2.94 $2.30$ 0.16 99.04 0.045 3.51
Table 2: Ablation study for the motion module and FusionTVO of our method during video face swapping.
@@ 4.4 Ablation study
Table 3: Ablation study for the impact of Face Former and ReferenceNet on identity injection.
@@
Reliability of motion module. Despite our efforts to design precise conditions for generating swapped faces with identical motion to the target face, the intrinsic diversity of diffusion models leads to a noticeable variation. It will generate faces with the same motion but slightly different details, making video face swapping unavailable. Here, we evaluate the importance of motion module from three aspects: identity preservation, motion fidelity, and video quality. Table 2 shows that DynamicFace could achieve great temporal consistency and aesthetic quality with FusionTVO.
Effectiveness of identity injection modules. ReferenceNet and Face Former aim to enhance identity preservation through detailed-level and high-level feature injection, respectively. To ensure a fair comparison, we trained three models under identical settings to evaluate the impact of these identity modules. As demonstrated in Table 3, both components significantly improve identity extraction. Specifically, ReferenceNet extracts fine-grained textures and selectively injects discriminative facial features into the main denoising UNet, while Face Former leverages high-level semantic embeddings derived from a pretrained face recognition model to reinforce identity coherence.
5 Conclusion
@@
We present a novel method named DynamicFace which leveraged the powerful pretrained diffusion model with delicately disentangled facial conditions, achieving promising face swapping results in both image and video domains. Experiments show that the designed facial conditions could give precise and unique evidence on required information (e.g., shape, expression, pose, lighting, and background). Several efforts have been taken to adapt the diffusion model to face swapping, including Face Former, ReferenceNet, Mixture-of-Guiders. Extensive experiments demonstrate that our framework can achieve superior results compared to previous methods, with better controllability and scalability. With above fine-grained facial conditions and FusionTVO, our method can achieve consistent video face swapping. We hope our attempt can inspire future work to further explore the formulation of face-swapping to achieve better results.
Supplementary Materials
Appendix A Detailed Ablation Study on Facial Composable Conditions
To rigorously validate the necessity of each facial condition, we conduct additional ablation experiments exclusively on the FaceForensics++ (FF++) dataset. We systematically evaluated the impact of each composable facial condition by training the model while removing individual components on purely public dataset, ensuring reproducibility. Crucially, Table 4 validated that no single condition could be removed without significant performance degradation, confirming their complementary yet disentangled contributions. Figure 8 shows the role of disentangled conditions and their collective effect.
[…]
$w/o.$ landmark 55.45 1.66 3.46 2.59 0.47
$w/o.$ normal 56.32 1.93 3.13 2.96 0.43
$w/o.$ lighting 56.30 1.34 2.89 2.27 0.46
$w/o.$ bg 73.05 2.43 4.01 2.66 0.48
Ours 56.22 1.46 2.60 2.06 0.33
Table 4: Ablation study on different facial conditions.
Appendix B Why Disentanglement Matters
The core strength of DynamicFace lies in its disentangled facial conditions, which enable independent control over identity, motion, and environmental attributes. We explore more possible applications to validate this:
Enhancing shape similarity of ID Preserving Text-to-Image Generation. Despite advancements in identity-preserving methods, existing approaches may still exhibit suboptimal ID similarity and inadequate face shape control. Shape-aware normal maps of DynamicFace anchor facial geometry to the source identity, enabling localized repairs (e.g. jaw realignment) without distorting the target’s expression or pose. We first generate portraits in different styles using ID Preserving methods and then enhance the generated results. Our method resolves identity leakage in the jaw and eye shape as shown in Figure 7.
Figure 7: Enhance results genearted by ID-Preserving works.
Motion-Consistent Artifact Restoring. By decoupling shape, expression, background, and illumination, DynamicFace could selectively restore corrupted regions without altering unrelated attributes (e.g., fixing a torn texture while preserving eye blinks). Our disentangled facial conditions enable precise control over motion attributes (e.g., expression, pose) while preserving the identity of the source face. This allows us to address facial artifacts in the results from other portrait generation works . Expression-related landmarks ensure repaired regions align with the original motion dynamics as shown in Figure 9, 10.
Figure 8: Impact of disentangled facial conditions on face swapping.
Figure 9: Disentangled priors enable motion-consistent facial artifact correction in generated videos
Figure 10: Disentangled priors enable motion-consistent facial artifact correction in generated videos
Appendix C Extended Real World Visualizations
We provide additional visual results on real-world videos to demonstrate robustness under challenging scenarios: extreme expressions and occlusions. The videos are sorted out together in the supplementary materials, including visual comparison with latest video face swapping methods.
Figure 11: Celebrity face swapping results under varied illumination and pose conditions
Figure 12: Celebrity face swapping results under varied illumination and pose conditions